If A baker is making fudge and his recipe calls for thirty 75% cocoa chocolate squares. He looks in the cupboard and discovers he only has 60% cocoa and 85% cocoa. How many of each should he use to make his fudge?
Answer by jorel1380(3719)

x+y=1
.60x+.85y=.75
60x+85y=75
60x+60y=60
25y=15
y=15/25=3/5
--------------
85x+85y=85
60x+85y=75
25x=10
x=10/25=2/5
The baker needs to use 2 parts of 60% cocoa for every 3 parts of 85% cocoa in his recipe..